OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 5x − 3y − z if x = −2, y = 2, and z = −3

OpenStudy (anonymous):

well remember PEMDAS and start with...?

jhonyy9 (jhonyy9):

5x-3y-z will be 5(-2)-3(2)-(-3)=-10-6+3=-16+3=-13